OpenStudy (caitlinnr14):

The sum of three consecutive integers is 267. What is the largest integer? Numerical Answers Expected!

OpenStudy (caitlinnr14):

@JA1 @SAValkyrie @Dboy21

OpenStudy (anonymous):

How can you show three consecutive integers using general terms?

OpenStudy (caitlinnr14):

....

OpenStudy (anonymous):

Im not allowed to give out answers for free anymore I got warned so I need to walk you through step by step

OpenStudy (anonymous):

If integers are consecutive, whats their difference?

OpenStudy (caitlinnr14):

ok i really dont understand and i need to learn this so thx u

OpenStudy (anonymous):

If there are three consecutive integers, you can show them as x-1, x, x+1. This is because the numbers are different by one. Do you understand what I mean?

OpenStudy (caitlinnr14):

oh yea!

OpenStudy (anonymous):

so by doing that, you can show this question using only x so it will look like this x-1+x+x+1=267

OpenStudy (caitlinnr14):

ok so basically i need to find 3 numbers that equal 267

OpenStudy (caitlinnr14):

?

OpenStudy (anonymous):

yeah so by using that idea of showing numbers into x-1, x, and x+1 you can set up an equation like I have shown above. By doing so, you can find the value of x

OpenStudy (caitlinnr14):

wait what?

OpenStudy (caitlinnr14):

@SAValkyrie

OpenStudy (anonymous):

Okay let's go back to the beginning. I said you can show the three consecutive integers by using this : x-1, x, x+1 correct?

OpenStudy (caitlinnr14):

yes

OpenStudy (anonymous):

The question above states: Sum of three consecutive integers that equals 267. So you need to find the sum(total) of the three integers. Since it's hard to do this by guess and check, we need to take a indirect method.

OpenStudy (caitlinnr14):

ok

OpenStudy (anonymous):

The three consecutive integers are x-1, x, x+1. Find the sum of those three terms and they should equal to 267 like this (x-1)+x+(x+1)=267

OpenStudy (caitlinnr14):

ok so like 184-90+183?

OpenStudy (caitlinnr14):

@SAValkyrie

OpenStudy (anonymous):

umm no.. you are supposed to find three integers that are consecutive so like this 88,89,90

OpenStudy (anonymous):

(x-1)+x+(x+1)=267 solve this equation, solve for x and you should get the answer for x which will give you the answer for this problem.

OpenStudy (caitlinnr14):

oh so lik 60,30,177?

OpenStudy (caitlinnr14):

@SAValkyrie

OpenStudy (anonymous):

eh no

OpenStudy (anonymous):

(x-1)+x+(x+1)=267 x+x+x-1+1=267 3x=267 x=89

OpenStudy (caitlinnr14):

ohhhhh!

OpenStudy (anonymous):

So the three numbers would be 88,89 and 90